OpenStudy (anonymous):

Polynomials! Can somebody please explain the steps to create a polynomial equation from roots -6, -1, 4?

OpenStudy (anonymous):

@phi Can you help me?

OpenStudy (phi):

the "roots" means x= -6 x= -1 or x= 4 rewrite each of those equations so all the terms are on the left side. can you do that ?

OpenStudy (anonymous):

x+6=0 x+1=0 x-4=0

OpenStudy (phi):

now multiply the equations like this: (x+6)(x+1)(x-4) = 0 that is your polynomial you can re-write it in "standard form" by multiplying it out

OpenStudy (anonymous):

so it would be x^3+3x^2-22x-24